Consider the real exchange rate. Suppose that the nominal exchange rate is 30, which is 10% lower than the level predicted by the absolute PPP. Answer how much the real exchange rate is.
1
Expert's answer
2020-03-18T14:02:40-0400

"Er=En\times\frac{P_f}{P_d}" = 30*0.9=27

Er - real exchange rate

En - nominal exchange rate

Pf/Pd - absolute PPP